OpenStudy (anonymous):

]Simplify by combining like terms. seventy-five times m minus thirty-six times m minus thirty-six plus seventy-five times m m one hundred fifty times m minus seventy-two forty times m plus thirty-nine one hundred fourteen times m minus thirty-six

OpenStudy (ash2326):

I'll show you the first one, others you'll try:D

OpenStudy (ash2326):

We are given seventy-five times m minus thirty-six times m minus thirty-six plus seventy-five times m this is \[ 75 \times m -36 \times m-36 + 75 \times m\] We have \[75m-36m-36+75m\] Let's group the like terms \[75m-36m+75m-36\] Let's combine like terms, we get \[114m-36\] Now you try the others